# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 3.4.1. Log smooth deformation theory
Recall that on a scheme $X$, a log structure is given by a sheaf of commutative monoids $\mathcal{M}_{X}$, together with a homomorphism of sheaves $\alpha_{X}:\mathcal{M}_{X}\to\mathcal{O}_{X}$ between multiplicative monoids, inducing an isomorphism $\alpha_{X}^{-1}(\mathcal{O}_{X}^{\times})\rightarrow\mathcal{O}_{X}^{\times}$. This isomorphism allows us to identify $\mathcal{O}_{X}^{\times}$ as a subsheaf of $\mathcal{M}_{X}$. We call a scheme $X$ endowed with a log structure a log scheme, and denote it by $X^{\dagger}:=(X,\mathcal{M}_{X})$ – we suppress the $\alpha_{X}$ from the notation, at situations it is clear what is is. Throughout this paper we assume that all log schemes are fine and saturated [79, I,§1.3]. The sheaf ¯$\overline{\mathcal{M}}_{X}:=\mathcal{M}_{X}/\mathcal{O}_{X}^{\times}$ known as the ghost sheaf or characteristic sheaf, captures the key combinatorial information about the log structure and in particular, the tropicalization $\Sigma(X)$ of $X$, which is an abstract polyhedral cone complex – see [1], §2.1 for details. In brief, the tropicalization $\Sigma(X)$ is given by a collection of cones along with face maps between them. For every geometric point $\bar{x}\rightarrow X$, there is one cone $\sigma_{\bar{x}}:=\mathrm{Hom}(\overline{\mathcal{M}}_{X,\bar{x}},\mathbb{R}_{\geq 0})$, and if $\bar{x}$ specializes to $\bar{y}$, there is a generization map ¯$\overline{\mathcal{M}}_{X,\bar{y}}\rightarrow\overline{\mathcal{M}}_{X,\bar{x}}$. This leads dually to a map $\sigma_{\bar{x}}\rightarrow\sigma_{\bar{y}}$, and the assumption that our log structures implies this is an inclusion of faces.
@@
An important log structure for us is the divisorial log structure, associated to a scheme $X$ with a divisor $D\subset X$. This log structure, denoted by $\mathcal{M}_{(X,D)}$, is given by the subsheaf of $\mathcal{O}_{X}$ consisting of functions invertible on the complement $X\setminus D$. In the situation when $X=\mathrm{Spec}\,\mathbb{C}[Q]$ is a toric variety, endowed with the divisorial log structure where $D:=\partial X$ given by the toric boundary divisor, the tropicalization $\Sigma(X)=(\mathbb{R}^{n},\Sigma)$ corresponds to the toric fan $\Sigma\subset\mathbb{R}^{n}$ for $X$.
@@
Given a log scheme $Y^{\dagger}=(Y,\mathcal{M}_{Y})$ and a morphism of schemes $f:X\to Y$, one can define an induced log structure, also known as pull-back log structure on $X$, which we denote by $f^{*}\mathcal{M}_{Y}$ – see [13, Defn.2.16]. For a monoid $Q$ such that $Q^{\times}=\{0\}$, pulling-back the divisorial log structure on the affine toric variety $X=\mathrm{Spec}\,\mathbb{C}[Q]$ to its torus fixed point $\mathrm{Spec}\,\mathbb{C}$, we obtain the standard log structure on $\mathrm{Spec}\,\mathbb{C}$ given by
@@
$Q\oplus\mathbb{C}^{\times}\longrightarrow\mathbb{C},\quad(q,a)\longmapsto\begin{cases}a,&q=0\\ 0,&q\neq 0.\end{cases}\,.$
Throughout this paper, unless stated otherwise, we assume $Q:=\mathbb{N}$ and denote the standard log point by ${\mathrm{Spec}\,\mathbb{C}}^{\dagger}:=(\mathrm{Spec}\,\mathbb{C},\mathbb{N}\oplus\mathbb{C}^{\times})$.
@@
A morphism of log schemes $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$ consists of a scheme theoretic morphism $f:X\rightarrow Y$, along with a map $f^{\flat}:f^{-1}\mathcal{M}_{Y}\rightarrow\mathcal{M}_{X}$ compatible with $f^{\#}:f^{-1}\mathcal{O}_{Y}\rightarrow\mathcal{O}_{X}$ via the structure homomorphisms $\alpha_{X}$ and $\alpha_{Y}$. Tropicalization of log schemes is functorial: given a morphism of log schemes $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$ inducing $f_{\mathrm{trop}}:\Sigma(X)\rightarrow\Sigma(Y)$, with a map of cones $\sigma_{\bar{x}}\rightarrow\sigma_{f(\bar{x})}$ induced by $\bar{f}^{\flat}:\overline{\mathcal{M}}_{Y,f(\bar{x})}\rightarrow\overline{\mathcal{M}}_{X,x}$. In the situation, we have a one-parameter degeneration of a scheme $X$, with central fiber $X_{0}$, the divisorial log structure $\mathcal{M}_{X,X_{0}}$ induces a log structure $\mathcal{M}_{X_{0}}$ on $X_{0}$. In this case, the tropicalization $\Sigma(X_{0})$ of the log scheme $(X_{0},\mathcal{M}_{X_{0}})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$ of $X_{0}$ – see for instance [15, Proposition 7.4.i].
@@
Similarly, as for scheme theoretic morphisms we have the notion of smoothness which can be defined using an infinitesimal lifting criterion, for log morphisms, we have the analog of this notion, given by log smoothness. We call a morphism log smooth, if and only if it is locally toroidal, in the sense of [57, (3.5)]. In the following sections, we study the deformation theory of the normal crossing surface $\mathcal{X}_{0}$, which satisfies “log smoothness”, since by Theorem 3.21 it is $d$-semistable, and as shown in [62] and in [57, Theorem 11.2], this implies log smoothness. While log smooth deformation theory is developed in a general set-up by Fumiharu Kato [57], it suffices for our purposes to consider the normal crossing situation, following [62]. We do so in the following section.
@@ 3.4.2. Smoothing of $\mathcal{X}_{0}$
Throughout this section, for every log morphism $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$, between two log schemes $X$ and $Y$ we denote by $T_{X^{\dagger}/Y^{\dagger}}$ (resp. $\Omega_{X^{\dagger}/Y^{\dagger}}$) the associated relative log tangent (resp. cotangent) sheaf, which reduces to a vector bundle in the situation when the morphism $f^{\dagger}$ is log smooth [58, Proposition 3.10].
@@
${\pi_{0}}^{\dagger}\colon{\mathcal{X}_{0}}^{\dagger}=(\mathcal{X}_{0},\mathcal{M}_{\mathcal{X}_{0}})\longrightarrow\mathrm{Spec}\,\mathbb{C}^{\dagger}\,,$ (3.2)
@@
Since $\mathcal{X}_{0}$ is $d$-semistable, by [62, Proposition 1.1] it admits a log structure which is log smooth over the standard log point. Since by construction $(P,\mathscr{P})$ is the dual intersection complex of $\mathcal{X}_{0}$, the statement about the tropicalization is immediate by the description of the tropicalization of a map, as explained in §3.4.1. ∎
The key ingredient to show the existence of the smoothing of $\mathcal{X}_{0}$, is the “vanishing result” shown in Lemma 3.25. Its proof uses the following observation for smooth log Calabi–Yau surfaces.
@@
Let $(X,D)$ be a smooth log Calabi–Yau surface, and denote by $D=D_{1}+\ldots+D_{n}$ the cyclically ordered irreducible components of $D$. Then, the components $\{D_{3},\ldots,,D_{n}\}$ are linearly independent in $H^{2}(X,\mathbb{C})$.
@@
We will show that, if there exists $a_{i}\in\mathbb{C}$ such that
$a_{3}D_{3}+\ldots a_{n}D_{n}=0$ (3.3)
then $a_{i}=0$ for all $3\leq i\leq n$. First, taking the intersection with $D_{2}$ of both sides of Equation (3.3), we get $a_{3}(D_{2}\cdot D_{3})+\ldots a_{n}(D_{2}\cdot D_{n})=0$. Since $D_{2}\cdot D_{3}=1$ and $D_{2}\cdot D_{i}=0$ for all $3\leq i\leq n$, as $D_{1}+\ldots+D_{n}$ is cyclically ordered, this gives us $a_{3}=0$. Thus, Equation (3.3) reduces to $a_{4}D_{4}+\ldots a_{n}D_{n}=0$. Now, taking the intersection with $D_{3}$ of both sides, analogously as before, we get $a_{4}=0$. Proceeding similarly, eventually we get $a_{i}=0$ for all $3\leq i\leq n$. ∎
The following Lemma 3.24 will also be crucial for the proof of the vanishing result, which follows afterwards. First, define
$\mathcal{Z}:=\bigcup_{\begin{subarray}{c}v\in\mathscr{P}^{[0]}\\ v\notin\partial P\end{subarray}}X^{v}$ (3.4)
the union of the proper irreducible components of $\mathcal{X}_{0}$, and
$\partial\mathcal{Z}:=\mathcal{Z}\cap\bigcup_{v\in(\partial P)_{\mathbb{Z}}}X^{v}\subset\mathcal{Z}$
the Cartier divisor on $\mathcal{Z}$ given by the union of curves in the double locus of $\mathcal{X}^{0}_{\mathscr{P}}$ which are in the intersection of a proper irreducible component with a non-proper irreducible component.
@@
We have $H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))=0.$
@@
We adapt to our situation the argument given in the proof of [33, Lemma 5.9] (where the log cotangent bundle is denoted by $\Lambda^{1}_{X}$) in the context of degenerations of K3 surfaces – see also [84, Theorem 5.2] for a similar result in the context of degenerations of Enriques surfaces.
As in [33, Corollary 3.6] or [62, Proof of Theorem 4.2], it follows from the definition of the log cotangent bundle that we have a short exact sequence
$0\rightarrow\Omega_{\mathcal{X}_{0}}/\tau_{\mathcal{X}_{0}}\rightarrow\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}\rightarrow F\rightarrow 0\,,$
where $\Omega_{\mathcal{X}_{0}}$ is the usual cotangent sheaf of $\mathcal{X}_{0}$, $\tau_{\mathcal{X}_{0}}$ is the torsion subsheaf of $\Omega_{\mathcal{X}_{0}}$, and the cokernel $F$ sits via the Poincaré residue map into the short exact sequence
$0\rightarrow\mathcal{O}_{\mathcal{X}_{0}}\rightarrow\bigoplus_{v\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{v}}\rightarrow F\rightarrow 0\,,$
induced by the decomposition $\mathcal{X}_{0}=\bigcup_{v\in\mathscr{P}^{[1]}}X^{v}$ into irreducible components. Using the exact sequence
$0\rightarrow\mathcal{O}_{0}\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}\mathcal{O}_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}\mathcal{O}_{X^{f}}\rightarrow 0\,,$
we rewrite $F$ as
$F=\mathrm{Ker}\left(\bigoplus_{e\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}\mathcal{O}_{X^{f}}\right)\,.$
In restriction to $\mathcal{Z}$, we obtain the short exact sequence
$0\rightarrow\Omega_{\mathcal{X}_{0}}|_{\mathcal{Z}}/\tau_{\mathcal{Z}}\rightarrow\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{Z}\rightarrow G\rightarrow 0\,,$ (3.5)
where $\tau_{\mathcal{Z}}$ is the torsion subsheaf of $\Omega_{\mathcal{X}_{0}}|_{\mathcal{Z}}$, and the cokernel $G$ sits via the Poincaré residue map into the short exact sequence
$0\rightarrow\mathcal{O}_{\mathcal{Z}}\rightarrow\bigoplus_{v\in\mathscr{P}_{\mathcal{Z}}^{[1]}}\mathcal{O}_{X^{v}}\oplus\bigoplus_{v\in(\partial P)_{\mathbb{Z}}}\mathcal{O}_{X^{v}\cap\mathcal{Z}}\rightarrow G\rightarrow 0\,,$
where $\mathscr{P}_{\mathcal{Z}}^{[1]}$ is the set of vertices $v\in\mathscr{P}^{[1]}$ such that $X^{v}_{\mathscr{P}}\subset\mathcal{Z}$, that is, $v\notin\partial P$. One can also describe $G$ as
$G=\mathrm{Ker}\left(\bigoplus_{e\in\mathscr{P}_{\mathcal{Z}}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}_{\mathcal{Z}}^{[2]}}\mathcal{O}_{X^{f}}\right)\,,$ (3.6)
where $\mathscr{P}_{Z}^{[1]}$ is the set of edges $e\in\mathscr{P}^{[1]}$ such that $X^{e}\subset\mathcal{Z}$, that is $e\not\subset\partial P$, and $\mathscr{P}_{\mathcal{Z}}^{[2]}$ is the set of faces $f\in\mathscr{P}^{[2]}$ whose three edges are in $\mathscr{P}_{\mathcal{Z}}^{[1]}$. We also denote by $\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}$ (resp. $\mathscr{P}^{[1]}_{\partial\mathcal{Z}}$) the set of edges $e\in\mathscr{P}^{[1]}$ such that $X^{e}\subset\mathcal{Z}$ and $X^{e}\not\subset\partial\mathcal{Z}$ (resp. $X^{e}\subset\mathcal{Z}$ and $X^{e}\subset\partial\mathcal{Z}$).
For every vertex $v\in\mathscr{P}^{[1]}_{\mathcal{Z}}$, the irreducible component $X^{v}_{\mathscr{P}}$ is a proper log Calabi–Yau surface, so $H^{0}(X^{v},\Omega)\subset H^{1}(X^{v},\mathbb{C})$ by the Hodge decomposition, and so we have $H^{0}(X^{v},\Omega)=0$ since $H^{1}(X^{v},\mathbb{C})=0$. On the other hand, for every $e\in\mathscr{P}^{[1]}_{\mathcal{Z}}$, we have $X^{e}\simeq\mathbb{P}^{1}$, and so $H^{0}(X^{e},\Omega_{X^{e}})=0$. Hence, the short exact sequence
$0\rightarrow\Omega_{\mathcal{X}_{\mathscr{P},0}}/\tau_{Z}\rightarrow\bigoplus_{v\in\mathscr{P}^{[1]}_{\mathcal{Z}}}\Omega_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}\Omega_{X^{e}}\rightarrow 0$
implies that we have $H^{0}(\mathcal{X}_{\mathscr{P},0},\Omega_{\mathcal{X}_{\mathscr{P},0}}/\tau_{Z})=0$ and
$H^{1}(\mathcal{X}_{\mathscr{P},0},\Omega_{\mathcal{X}_{\mathscr{P},0}}/\tau_{Z})$ $=\mathrm{Ker}\left(\bigoplus_{v\in\mathscr{P}^{[1]}_{\mathcal{Z}}}H^{1}(X^{v},\Omega_{X^{v}})\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}H^{1}(X^{e},\Omega_{X^{e}})\right)$
$=\mathrm{Ker}\left(\bigoplus_{v\in\mathscr{P}^{[1]}_{\mathcal{Z}}}H^{2}(X^{v},\mathbb{C})\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}H^{2}(X^{e},\mathbb{C})\right)\,.$
Thus, the long exact sequence in cohomology associated to the short exact sequence (3.5) takes the form
$0\rightarrow H^{0}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})\rightarrow H^{0}(\mathcal{X}_{0},G)\rightarrow\bigoplus_{v\in\mathscr{P}^{[1]}_{\mathcal{Z}}}H^{2}(X^{v},\mathbb{C})\,.$
It follows from the exact sequence in Equation (3.5) and from the description of $G$ in Equation (3.6) that we have an exact sequence
[…]
where
$H=\mathrm{Ker}\left(\bigoplus_{e\in\mathscr{P}_{\mathcal{Z}\setminus\partial\mathcal{Z}}^{[1]}}H^{0}(X^{e},\mathcal{O}_{X^{e}})\rightarrow\bigoplus_{f\in\mathscr{P}_{\mathcal{Z}}^{[2]}}H^{0}(X^{f},\mathcal{O}_{X^{f}})\right)\,,$
and the map $\delta$ in (3.7) is induced by the maps
$H^{0}(X^{e},\mathcal{O}_{X^{e}})\simeq\mathbb{C}$ $\longrightarrow H^{2}(X^{v},\mathbb{C})$
11 11 $\longmapsto[X^{e}]$
for every $v$ and $e$ such that $e$ is adjacent to $v$, that is, such that $X^{e}\subset X^{v}$.
To prove Lemma 3.24, we have to show that $H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))=0$, that is, by (3.7), that the map $\delta$ is injective. Explicitly, an element of $H$ is a sequence $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ of numbers $a_{e}\in\mathbb{C}$, satisfying the balancing condition $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ each time we have three edges $e_{1},e_{2},e_{3}\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}$ bounding a triangle $f\in\mathscr{P}^{[2]}_{\mathcal{Z}}$. Let $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ be in the kernel of $\delta$, that is, that the image of $\sum_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}a_{e}[X^{e}]$ in $H^{2}(X^{v},\mathbb{C})$ is zero for every $v\in\mathscr{P}^{[1]}_{\mathcal{Z}}$. For every vertex $v\in\mathscr{P}^{[0]}$, define the distance to the boundary $d(v)$ as $d(v)=0$ if $v\in\partial P$, and else as the smallest integer such that there exists a sequence $(e_{1},\dots,e_{d(v)})$ of edges $e_{i}\in\mathscr{P}^{[1]}$ such that $e_{1}$ is adjacent to $v$, $e_{i+1}$ is adjacent to $e_{i}$, and $e_{d(v)}\in\partial P$. We will prove that $a_{e}=0$ for all $e\in\mathscr{P}^{[1]}_{\mathcal{Z}}$, by showing by induction on $d(v)\geq 0$ that, for every vertex $v\in\mathscr{P}^{[0]}$, we have $a_{e}=0$ for all edges $e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}$ adjacent to $v$.
For the initialization of the induction, note that if $v\in\mathscr{P}^{[1]}$ is such that $d(v)=0$, then $v\in\partial P$, so we have $e\notin\mathscr{P}^{1}_{\mathcal{Z}\setminus\partial\mathcal{Z}}$ for every edge $e$ adjacent to $v$, and so the result is empty is this case. For the induction step, let $d\in\mathbb{Z}_{\geq 1}$, and assume that the result holds for all vertices $v^{\prime}$ with $d(v^{\prime})<d$. Let $v\in\mathscr{P}^{[0]}_{\mathcal{Z}}$ such that $d(v)=d$. Then, there exists an edge $e_{1}$ adjacent to $v$ which is also adjacent to a vertex $v^{\prime}\in\mathscr{P}^{[1]}$ with $d(v^{\prime})<d$. Let $f\in\mathscr{P}^{[2]}$ be a triangle with side $e_{1}$, and denote by $e_{2}$ the other side adjacent to $v$, and by $e_{3}$ the other side adjacent to $v^{\prime}$. By the induction hypothesis, we have $a_{e_{1}}=a_{e_{3}}=0$, and so it follows from the balancing condition $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ that we also have $a_{e_{2}}=0$. Therefore, the restriction of $\sum_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}a_{e}[X^{e}]$ to $X^{v}$ is a linear combination of the boundary divisors of $X^{v}$ not containing the consecutive components $X^{e_{1}}$ and $X^{e_{2}}$. Hence, by Lemma 3.23, the condition that $\sum_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}a_{e}[X^{e}]$ is zero in $H^{2}(X^{v},\mathbb{C})$ implies that $a_{e}=0$ for all edges $e$ adjacent to $v$. This concludes the proof that $\delta$ is injective, and so the proof of Lemma 3.24. ∎
Now, we are ready to prove the vanishing result.
@@
We have $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}})=0\,$ where $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is the log tangent bundle of $\mathcal{X}_{0}^{\dagger}$ over Specnormal-$\mathrm{Spec}\,\mathbb{C}^{\dagger}$.

### v3

@@ 3.3.1. Log smooth deformation theory
Recall that on a scheme $X$, a log structure consists of a sheaf of commutative monoids $\mathcal{M}_{X}$ and a homomorphism of sheaves $\alpha_{X}:\mathcal{M}_{X}\to\mathcal{O}_{X}$ between multiplicative monoids, inducing an isomorphism $\alpha_{X}^{-1}(\mathcal{O}_{X}^{\times})\rightarrow\mathcal{O}_{X}^{\times}$. This isomorphism allows us to identify $\mathcal{O}_{X}^{\times}$ as a subsheaf of $\mathcal{M}_{X}$. A scheme $X$ endowed with a log structure is called a log scheme, and is denoted by $X^{\dagger}:=(X,\mathcal{M}_{X})$ – we suppress $\alpha_{X}$ from the notation when its meaning is clear from context. Throughout this paper we assume that all log schemes are fine and saturated [84, I,§1.3].
The sheaf $\overline{\mathcal{M}}_{X}:=\mathcal{M}_{X}/\mathcal{O}_{X}^{\times}$, known as the ghost sheaf or characteristic sheaf, encodes the key combinatorial aspects of the log structure and, in particular, determines the tropicalization $\Sigma(X)$ of $X$, which is an abstract polyhedral cone complex – see [1], §2.1 for details. In brief, the tropicalization $\Sigma(X)$ consists of a collection of cones along with face maps between them. For every geometric point $\bar{x}\rightarrow X$, there is an associated cone $\sigma_{\bar{x}}:=\mathrm{Hom}(\overline{\mathcal{M}}_{X,\bar{x}},\mathbb{R}_{\geq 0})$. Moreover, if $\bar{x}$ specializes to $\bar{y}$, there is a generization map $\overline{\mathcal{M}}_{X,\bar{y}}\rightarrow\overline{\mathcal{M}}_{X,\bar{x}}$, which dually induces a map $\sigma_{\bar{x}}\rightarrow\sigma_{\bar{y}}$.
@@
An important log structure for us is the divisorial log structure, which is associated with a scheme $X$ with a divisor $D\subset X$. This log structure, denoted by $\mathcal{M}_{(X,D)}$, is defined as the subsheaf of $\mathcal{O}_{X}$ consisting of functions that are invertible on the complement $X\setminus D$. In the situation when $X=\mathrm{Spec}\,\mathbb{C}[Q]$ is a toric variety endowed with the divisorial log structure induced by the toric boundary divisor $D:=\partial X$, the tropicalization $\Sigma(X)=(\mathbb{R}^{n},\Sigma)$ corresponds to the toric fan $\Sigma\subset\mathbb{R}^{n}$ for $X$.
@@
Given a log scheme $Y^{\dagger}=(Y,\mathcal{M}_{Y})$ and a morphism of schemes $f:X\to Y$, one can define an induced log structure, also known as pull-back log structure on $X$, which we denote by $f^{*}\mathcal{M}_{Y}$ – see [13, Definition 2.16]. For a monoid $Q$ such that $Q^{\times}=\{0\}$, pulling back the divisorial log structure on the affine toric variety $X=\mathrm{Spec}\,\mathbb{C}[Q]$ to its torus fixed point $\mathrm{Spec}\,\mathbb{C}$, defines a log structure on $\mathrm{Spec}\,\mathbb{C}$ given by
@@
$Q\oplus\mathbb{C}^{\times}\longrightarrow\mathbb{C},\quad(q,a)\longmapsto\begin{cases}a,&q=0\\ 0,&q\neq 0\,.\end{cases}$ (3.2)
[…]
@@
A morphism of log schemes $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$ consists of a scheme theoretic morphism $f:X\rightarrow Y$, along with a map $f^{\flat}:f^{-1}\mathcal{M}_{Y}\rightarrow\mathcal{M}_{X}$ that is compatible with $f^{\#}:f^{-1}\mathcal{O}_{Y}\rightarrow\mathcal{O}_{X}$ via the structure homomorphisms $\alpha_{X}$ and $\alpha_{Y}$. The tropicalization of log schemes is functorial: a morphism of log schemes $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$ induces a map of cone complexes $f_{\mathrm{trop}}:\Sigma(X)\rightarrow\Sigma(Y)$. Given a one-parameter degeneration of a scheme $X$, with reduced central fiber $X_{0}$, the divisorial log structure $\mathcal{M}_{(X,X_{0})}$ induces a log structure $\mathcal{M}_{X_{0}}$ on $X_{0}$. The tropicalization $\Sigma(X_{0})$ of the log scheme $(X_{0},\mathcal{M}_{X_{0}})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$ of $X_{0}$ – see for instance [15, Proposition 7.4.i].
@@
Similarly to scheme-theoretic morphisms, for which there is the notion of smoothness defined using the infinitesimal lifting criterion, there is for log morphisms an analogous notion known as log smoothness. In the following sections, we employ log smooth deformation theory to study $d$-semistable open Kulikov surfaces. While log smooth deformation theory is developed in a general framework in [63], it suffices for our purposes to consider the normal crossing situation, following [68].
@@ 3.3.2. Smoothing of $\mathcal{X}_{0}$
Throughout this section, for every log morphism $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$ between two log schemes $X$ and $Y$, we denote by $T_{X^{\dagger}/Y^{\dagger}}$ (resp. $\Omega_{X^{\dagger}/Y^{\dagger}}$) the associated relative log tangent (resp. cotangent) sheaf, which reduces to a vector bundle in the situation when the morphism $f^{\dagger}$ is log smooth [64, Proposition 3.10].
@@
${\pi_{0}}^{\dagger}\colon{\mathcal{X}_{0}}^{\dagger}=(\mathcal{X}_{0},\mathcal{M}_{\mathcal{X}_{0}})\longrightarrow\mathrm{pt}_{\mathbb{N}}\,$, (3.3)
@@
Since $\mathcal{X}_{0}$ is $d$-semistable, by [68, Proposition 1.1] it admits a log structure which is log smooth over the standard log point $\mathrm{pt}_{\mathbb{N}}$. Since by construction $(P,\mathscr{P})$ is the dual intersection complex of $\mathcal{X}_{0}$, the statement about the tropicalization is immediate by the description of the tropicalization of a map, as explained in §3.3.1. ∎
The key ingredient to show the existence of the smoothing of $\mathcal{X}_{0}$, is the “vanishing result” shown in Lemma 3.18. Its proof uses the following observations on smooth log Calabi–Yau surfaces and open log Calabi–Yau surfaces.
@@
Let $(X,D)$ be a smooth log Calabi–Yau surface, and denote by $D=D_{1}+\ldots+D_{n}$ the cyclically ordered irreducible components of $D$. Then, the components $D_{3},\ldots,,D_{n}$ are linearly independent in $H^{2}(X,\mathbb{C})$.
@@
Let $a_{3}D_{3}+\ldots a_{n}D_{n}=0$ be a linear relation between $D_{3},\dots,D_{n}$ with coefficients $a_{i}\in\mathbb{C}$. Taking the intersection with $D_{2}$, we obtain $a_{3}(D_{2}\cdot D_{3})+\ldots+a_{n}(D_{2}\cdot D_{n})=0$. Since $D_{1}+\ldots+D_{n}$ is cyclically ordered, we have $D_{2}\cdot D_{3}=1$ and $D_{2}\cdot D_{i}=0$ for all $3\leq i\leq n$, and so $a_{3}=0$. Thus, the linear relation reduces to $a_{4}D_{4}+\ldots+a_{n}D_{n}=0$. Taking the intersection with $D_{4}$, we similarly deduce that $a_{4}=0$. Iterating the argument, we eventually get $a_{i}=0$ for all $3\leq i\leq n$. ∎
@@
Let $(X,D)$ be an open log Calabi–Yau surface, and denote by $D=D_{0}+D_{1}+\ldots+D_{n-1}+D_{n}$ the ordered irreducible components of $D$, so that $D_{0}$ and $D_{n}$ are non-compact, and $D_{1},\cdots,D_{n-1}$ are compact. Then, the compact components $D_{1},\ldots,,D_{n-1}$ are linearly independent in the cohomology group with compact support $H^{2}_{c}(X,\mathbb{C})$.
@@
Analogue to the proof of Lemma 3.16, by successively intersecting a relation $\sum_{i=1}^{n-1}a_{i}D_{i}=0$ with $D_{0}$, $D_{1}$, …, $D_{n-1}$. ∎
@@
Let $\mathcal{X}_{0}$ be a $d$-semistable open Kulikov surface. Then, we have
$H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}})=0\,$,
where $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}}$ is the log tangent bundle of $\mathcal{X}_{0}^{\dagger}$ over the standard log point $\mathrm{pt}_{\mathbb{N}}$.